OpenStudy (anonymous):

I need a well thought through answer to this, but any answers are appreciated. Which method for systems of equations (Graphing, Elimination, Substitution) do you prefer, and why?

OpenStudy (anonymous):

I like substitution and elimination. Not sure I learned graphing though. Well it's easier and fun to do.

OpenStudy (anonymous):

Thank you, Natshane.

OpenStudy (anonymous):

you're welcome :) Basically to me, substitution and elimination is the basic to solving any expressions.

OpenStudy (anonymous):

Very true. I was asked by my teacher to pick which one of the three that I prefered. I chose graphing and told her it was because it was easiest to me, but she didn't take it as an answer. Any suggestions?

OpenStudy (anonymous):

The absolute best method is matrix elimination. This way you can solve a system of \(n\) equations with \(n\) unknowns in a minute.

OpenStudy (anonymous):

@badreferences I have not learned the 'matrix' elimination. And I don't know what that little 'n' like figure is.

OpenStudy (anonymous):

I'm sorry I'm not sure if I learned graphing, haha, or maybe I can't recall. If you want to choose elimination, maybe you can say that it's easier to eliminate the same unknown to find the another unknown? And then using that value of the unknown and substitute it in the first equation to find the other unknown? lol, it sounds confusing, but here's an example. 2a + b = 6 a - b = 3 plus both of the equations (elimination of b ) and you get 3a = 9, a = 3. Then substitute a = 3 into 2a + b = 6 and you get b = 0

OpenStudy (anonymous):

Oh, alright. That makes alot more sense. Thank you very much, Nat. ^_^

OpenStudy (anonymous):

you're welcome! I hope I do help:)

OpenStudy (anonymous):

I'll tell you what the teacher says later XD (:

OpenStudy (anonymous):

Haha! I'm curious, keep me posted :)